Juan made a 2L of lemonade that consists of 40% pure lemon juice.How much water must he add to dilute it to 30% lemon juice?
Answer by mananth(16946)

Juan made a 2L of lemonade that consists of 40% pure lemon juice.How much water must he add to dilute it to 30% lemon juice?
...
water ---- 0% juice x liters
lemonade 40% juice. 2L
Total = x+2 liters ------ 30 %
..
0+0.4*2=(x+2)*0.3
0.8=0.3x+0.6
0.8-0.6=0.3x
0.2=0.3x
/0.3
x=0.2/0.3
x=0.66 liters of water to be added.
...
m.ananth@hotmail.ca